OpenStudy (anonymous):

In standard circle form 3x^2+3y^2x+6y+10/3=48

OpenStudy (anonymous):

Someone help me

OpenStudy (anonymous):

Anyone :(

OpenStudy (anonymous):

9x^2+9y^2+18y=144-10 which is 9x^2+9y^2+18y=134

OpenStudy (anonymous):

compare it with x^2-2xa+a^2+y^2-2by+b^2=r^2

OpenStudy (anonymous):

equation of circle centred at a,b